OpenStudy (goten77):

in calculus we cannot solver for a anti derivative of e^cos(x) but is there any other impossible anti derivatives?

OpenStudy (anonymous):

\[\int_0^xe^{\cos(t)}dt\] would be an anti derivative

OpenStudy (cruffo):

nothin's "impossible" :) but you may not be able to find a closed for for \[\int f(x) \text{d}x\]

OpenStudy (cruffo):

Sorry, the first for should be form

OpenStudy (goten77):

XD i dont think many more exist except like 5

OpenStudy (zzr0ck3r):

5 what?

OpenStudy (cruffo):

Would you consider \[\huge \int e^{x^2} \text{d} x\] to be integrable by regular methods (i.e. substitution, by-parts,...)?